# Semiformal method I. Pólya’s theorem for the complete graph $K_{\mathbb{N}}$ with complex edge weights

### v3

@@ 4 Extensions of Pólya’s theorem with $v\neq 1$
The case when the vertex visited by walks differs from their starting vertex is more complicated. The grid graph $\mathbb{Z}^{d}$ is vertex-transitive, but so far, this has been irrelevant. For $\overline{v}\neq\overline{0}$ vertex-transitivity becomes important, and we generalize it to $K_{\mathbb{N}}$ as follows. Let $v\in\mathbb{N}\setminus\{1\}$ with $v\neq 1$. We say that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $v$-transitive if there is a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that $f(1)=v$ and $h(f[e])=h(e)$ for every $e\in\mathbb{N}_{2}$.
We again introduce some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ are the walks visiting the vertex $v$ ($\neq 1$); necessarily in a step $i>0$. We are interested in the generating function
@@
The generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$ are as before. We define
$B_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{b,v}(n))x^{n}\,,}$
$C_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{c,v}(n))x^{n}\,\text{ and}}$
$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e,v}(n))x^{n}\,,}$
all in $\mathbb{C}[[x]]$. Here $W_{b,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$. $W_{c,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding $v$. Finally, $W_{e,v}(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding both $1$ and $v$. We have $a_{0}^{h,v}=c_{0}^{h,v}=c_{0}^{h}=e_{0}^{h,v}=0$ and $b_{0}^{h}=b_{0}^{h,v}=d_{0}^{h}=1$.
We obtain an analog of Proposition 3.1. The proof is similar. We omit details and argue only on the high level of semiformal method.
@@
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and generating functions $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ be as above. The following relations hold between them.
@@
Proof. 1. Every walk $w$ weight-counted by $A_{h,v}(x)$ splits at the first visit of $v$ as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $C_{h,v}(x)$ and $w_{2}$ are arbitrary walks starting at $v$. By the $v$-transitivity of $h$, walks $w_{2}$ are weight-counted by $D_{h}(x)$.
2. It suffices to prove the first equality. $B_{h}(x)$ weight-counts walks $w$ starting and ending at $1$. Those avoiding $v$ are weight-counted by $B_{h,v}(x)$. If $w$ visits $v$, then it uniquely splits at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$. Walks $w_{1}$ start at $1$, end at $v$ and between avoid $v$. Walks $w_{2}$ start and end at $v$. Walks $w_{3}$ start at $v$, end at $1$ and between avoid $v$. We reverse walks $w_{3}$ and see that both $w_{1}$ and $w_{3}$ are weight-counted by $C_{h,v}(x)$. Due to the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$.
@@
For $w=0$ or $z=0$ it equals $\{0\}$. Else it has two non-zero elements differing by sign. If $w=1$, we write just $\mathrm{sqrt}(z)$.
@@
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. Suppose that $D_{h}(1)_{1}$, $C_{h,v}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h,v}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
@@
Proof. By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By Proposition 2.19, $A_{h,v}(1)_{1}=C_{h,v}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By item 2 of Proposition 3.1 and by Proposition 2.20, $B_{h}(1)_{1}\neq 0$. Thus by Proposition 2.22, $C_{h,v}(1)_{1}^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
We assume the existence of $E_{h,v}(1)_{1}$ instead of $C_{h}(1)_{1}$ and get the following variant of Theorem 4.2.
@@
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. Suppose that $E_{h,v}(1)_{1}$, $D_{h}(1)_{1}$, $C_{h,v}(1)_{1}$, $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$ and $A_{h,v}(1)_{1}$ exist. Then
@@
where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret $\frac{0}{0}$ as $1$, so that $A_{h,v}(1)_{1}=0$.
Proof. For $B_{h}(1)_{1}\neq 0$ we use the previous proof. If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.19 and 2.18 that also $B_{h,v}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.19 give $C_{h,v}(1)_{1}=0$. Finally, item 1 of Proposition 4.1 and Proposition 2.19 give $A_{h,v}(1)_{1}=0$. $\Box$
@@
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight such that $v\in V(h)$. Suppose that $D_{h}(1)_{2}$ and $B_{h,v}(1)_{2}$ exist, and that $c=\sum_{n\geq 1}|b_{n}^{h}|<1$. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. Let $B_{h,v}(x)\neq B_{h}(x)$ and $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1$}
then $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $A_{h,v}(1)_{2}$ exists and
@@
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}$ exists, $B_{h}(1)_{2}\neq 0$, $C_{h,v}(1)_{2}$ exists and $(C_{h,v}(1)_{2})^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}$. By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By item 1 of Proposition 2.25, $A_{h,v}(1)_{2}$ exists and $A_{h,v}(1)_{2}=C_{h,v}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
Theorem 4.5 (9: $v\neq 1$, gen, $U(1)_{1}=+\infty$)
Let $v\in\mathbb{N}\setminus\{1\}$ and let $h\colon\mathbb{N}_{2}\to\mathbb{R}_{\geq 0}$ be a $v$-transitive light weight such that $v\in V(h)$. If $D_{h}(1)_{1}=+\infty$ then
$A_{h,v}(1)_{1}=+\infty\,.$
Proof. Let $m\in\mathbb{N}$ be the minimum length of a walk in $G(h)$ joining $1$ and $v$. The inner vertices of this shortest walk differ from $v$ and we get that $c^{h,v}_{m}>0$. Since by item 1 of Proposition 4.1 we have $a_{n}^{h,v}=\sum_{j=0}^{n}c_{j}^{h,v}d_{n-j}^{h}$, we conclude that $a_{n+m}^{h,v}\geq c_{m}^{h,v}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}a_{j+m}^{h,v}\geq c_{m}^{h,v}\cdot\sum_{j=0}^{n}d_{j}^{h}$}
for every $n\in\mathbb{N}_{0}$ and $A_{h,v}(1)_{1}=+\infty$. $\Box$
We proceed to convex weights.
@@
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $C_{h,v}(1)_{1}$, $C_{h}(1)_{1}$, $B_{h,v}(1)_{1}$ and $B_{h}(1)_{1}$ exist. Then $B_{h}(1)_{1}\neq 0$ and
$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\,.$
Proof. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-x}$. By Proposition 2.15, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}$. By item 2 of Proposition 3.1 and by Proposition 2.20, $B_{h}(1)_{1}\neq 0$. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.22, $C_{h,v}(1)_{1}^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
We again assume the existence of $E_{h,v}(1)_{1}$ instead of $C_{h}(1)_{1}$ and get the following variant of the theorem.
@@
Let $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive convex light weight such that $v\in V(h)$. Suppose that $E_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$, $B_{h,v}(1)_{1}$ and $B_{h}(1)_{1}$ exist. Then
$\lim_{n\to\infty}a_{n}^{h,v}\in{\textstyle\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}}\big)\ \ (\subset\mathbb{C})\,$,
where if $B_{h}(1)_{1}=0$ then $B_{h,v}(1)_{1}=0$ and we interpret $\frac{0}{0}$ as $1$, so that the limit is zero.
Proof. If $B_{h}(1)_{1}\neq 0$, the previous proof works. If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.19 and 2.18 that also $B_{h,v}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.19 give $C_{h,v}(1)_{1}=0$. Thus, as in the previous proof, $\lim_{n\to\infty}a_{n}^{h,v}=C_{h,v}(1)_{1}=0$. $\Box$

### v6

@@ 4 Extensions of Pólya’s theorem with $v\neq 1$
In Theorems 4.34.10 we establish the second group of generalizations of Pólya’s theorem, in the more complicated case when the vertex to be visited by walks differs from their starting vertex. The grid graph $\mathbb{Z}^{d}$ is vertex-transitive, which has so far been irrelevant. For $\overline{v}\neq\overline{0}$, vertex-transitivity becomes relevant. We generalize it to $K_{\mathbb{N}}$ as follows. Let $v\in\mathbb{N}\setminus\{1\}$. We say that a weight $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is $v$-transitive if there is a bijection $f\colon\mathbb{N}\to\mathbb{N}$ such that $f(1)=v$ and $h(f[e])=h(e)$ for every $e\in\mathbb{N}_{2}$.
We again introduce some generating functions. Let $n\in\mathbb{N}_{0}$, $v\in\mathbb{N}\setminus\{1\}$ and $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a light weight. Recall that $W(n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, and that $W(v,n)\subset W(n)$ is the subset of walks visiting the vertex $v$ ($\neq 1$); necessarily in a step $i>0$. We are interested in the generating function
@@
In the previous section we introduced the generating functions $B_{h}(x)$, $C_{h}(x)$ and $D_{h}(x)$. Now we define
$B_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}b_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{b}(v,\,n))x^{n}\,,}$
$C_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}c_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{c}(v,\,n))x^{n}\,\text{ and}}$
$E_{h,v}(x)$ $=$ ${\textstyle\sum_{n\geq 0}e_{n}^{h,v}x^{n}=\sum_{n\geq 0}h(W_{e}(v,\,n))x^{n}\,,}$
all in $\mathbb{C}[[x]]$. Here $W_{b}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting and ending at $1$ and avoiding $v$. $W_{c}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding $v$. Finally, $W_{e}(v,n)$ is the set of walks of length $n$ in $K_{\mathbb{N}}$ starting at $1$, ending at $v$ and between avoiding both $1$ and $v$. The constant coefficients are $a_{0}^{h,v}=c_{0}^{h,v}=c_{0}^{h}=e_{0}^{h,v}=0$ and $b_{0}^{h}=b_{0}^{h,v}=d_{0}^{h}=1$.
We obtain an analog of Proposition 3.1. Since the proof is similar, we omit details and argue only on the level of semiformal method.
@@
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ be a $v$-transitive light weight and $A_{h,v}(x)$, $B_{h}(x)$, $B_{h,v}(x)$, $C_{h,v}(x)$, $D_{h}(x)$ and $E_{h,v}(x)$ be the above generating functions. The following relations hold.
@@
Proof. 1. Walks $w$ weight-counted by $A_{h,v}(x)$ split at the first visit of $v$ as $w=w_{1}w_{2}$. Walks $w_{1}$ are weight-counted by $C_{h,v}(x)$ and $w_{2}$ are arbitrary walks starting at $v$. By the $v$-transitivity of $h$, walks $w_{2}$ are weight-counted by $D_{h}(x)$.
2. It suffices to prove the first equality. $B_{h}(x)$ weight-counts walks $w$ starting and ending at $1$. Those avoiding $v$ are weight-counted by $B_{h,v}(x)$. If $w$ visits $v$, it uniquely splits at the first and last visits of $v$ as $w=w_{1}w_{2}w_{3}$. Walks $w_{1}$ start at $1$, end at $v$ and between avoid $v$. Walks $w_{2}$ start and end at $v$. Walks $w_{3}$ start at $v$, end at $1$ and between avoid $v$. We reverse walks $w_{3}$ and see that both $w_{1}$ and $w_{3}$ are weight-counted by $C_{h,v}(x)$. Due to the $v$-transitivity of $h$, the middle walks $w_{2}$ are weight-counted by $B_{h}(x)$.
@@
For $w=0$ or $z=0$ it equals $\{0\}$. Else it has two non-zero elements differing only by sign. If $w=1$, we write just $\mathrm{sqrt}(z)$.
Before we turn to generalizations of Pólya’s theorem in the case $v\neq 1$, we utilize the square in item 2 of Proposition 4.1. If $h\colon X\to\mathbb{R}$ is an absolutely convergent real series and $Y\subset X$, then in general the sum $\sum_{x\in Y}h(x)$ of the subseries need not be smaller than the sum $\sum_{x\in X}h(x)$. However, due to the mentioned square, we obtain the next monotonicity result for the sums
$b_{n}^{h,v}=\sum_{w\in W_{b}(v,\,n)}h(w)\,\text{ and }\,b_{n}^{h}=\sum_{w\in W_{b}(n)}h(w)\,.$}
It is expressed in terms of the generating functions $B_{h,v}(x)=\sum_{n\geq 0}b_{n}^{h,v}x^{n}$ and $B_{h}(x)=\sum_{n\geq 0}b_{n}^{h}x^{n}$.
Theorem 4.2
Let $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{R}$ be a $v$-transitive light weight and let $B_{h,v},B_{h},C_{h,v}\in H_{1}$. Then for every real $c\in(-1,1)$ the following holds.
  1. 1.
    If $B_{h}(c)=0$ then also $B_{h,v}(c)=0$.
  2. 2.
    If $B_{h}(c)\neq 0$ then
    $\frac{B_{h,v}(c)}{B_{h}(c)}\leq 1\,.$
Proof. This follows from relations 2 in Proposition 4.1. $\Box$
We proceed to the remaining generalizations of Pólya’s theorem.
@@
If $v$ is in $\mathbb{N}\setminus\{1\}$, if $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$, and if $A_{h,v}(1)_{1}$, $B_{h}(1)_{1}\neq 0$, $B_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$ and $D_{h}(1)_{1}$ exist, then
@@
Proof. By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By Proposition 2.20, $A_{h,v}(1)_{1}=C_{h,v}(1)_{1}D_{h}(1)_{1}$. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. Thus by Proposition 2.22, $\big(C_{h,v}(1)_{1}\big)^{2}=1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}$. The stated formula follows. $\Box$
If we replace the assumption of non-vanishing of $B_{h}(1)_{1}$ with the existence of $E_{h,v}(1)_{1}$, we obtain the following variant of the theorem.
@@
If $v$ is in $\mathbb{N}\setminus\{1\}$, if $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$, and if $A_{h,v}(1)_{1}$, $B_{h}(1)_{1}$, $B_{h,v}(1)_{1}$, $C_{h,v}(1)_{1}$, $D_{h}(1)_{1}$ and $E_{h,v}(1)_{1}$ exist, then
@@
where for $B_{h}(1)_{1}=0$ we have $B_{h,v}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then $A_{h,v}(1)_{1}=0$.
Proof. If $B_{h}(1)_{1}\neq 0$, we use the previous proof. If $B_{h}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1, and Propositions 2.20 and 2.19 that also $B_{h,v}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.20 give $C_{h,v}(1)_{1}=0$. Finally, item 1 of Proposition 4.1 and Proposition 2.20 give $A_{h,v}(1)_{1}=0$. $\Box$
@@
Suppose that $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive light weight with $v\in V(h)$,
$c=\sum_{n\geq 1}|b_{n}^{h}|<1$}
and that $B_{h,v}(1)_{2}$ and $D_{h}(1)_{2}$ exist. If $B_{h,v}(x)=B_{h}(x)$ then $A_{h,v}(x)=0$. For $B_{h,v}(x)\neq B_{h}(x)$ let $k\in\mathbb{N}_{0}$ be minimum such that $b_{k}^{h,v}\neq b_{k}^{h}$. If
$\big(|b_{k}^{h,v}-b_{k}^{h}|\cdot(1-c)\big)^{-1}\sum_{n>k}|b_{n}^{h,v}-b_{n}^{h}|<1\,$,}
then $B_{h}(1)_{2}\neq 0$ and
@@
Let $B_{h,v}(x)\neq B_{h}(x)$, $k$ be as stated and let the displayed bound hold. By item 2 of Proposition 4.1 we have $C_{h,v}(x)^{2}=1-\frac{B_{h,v}(x)}{B_{h}(x)}$. By Proposition 2.30, $B_{h}(1)_{2}\neq 0$ and
$(C_{h,v}(1)_{2})^{2}=1-\frac{B_{h,v}(1)_{2}}{B_{h}(1)_{2}}\,.$}
By item 1 of Proposition 4.1 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)$. By item 1 of Proposition 2.24, $A_{h,v}(1)_{2}=C_{h,v}(1)_{2}D_{h}(1)_{2}$. The stated formula follows. $\Box$
Theorem 4.6 (9: $v\neq 1$, gen, $U(1)_{1}=\infty$)
If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\xi\mathbb{R}_{\geq 0}$ is a $v$-transitive light weight with $v\in V(h)$ and if $(D_{h})_{1/\xi}(1)_{1}=+\infty$, then
$(A_{h,v})_{1/\xi}(1)_{1}=+\infty\,.$
Proof. We denote the minimum length of a walk joining $1$ and $v$ in $G(h)$ by $m$. The inner vertices of it differ from $v$. By the subseries bound principle we get $\xi^{-m}c^{h,v}_{m}>0$. By item 1 of Proposition 4.1 we have $a_{n}^{h,v}=\sum_{j=0}^{n}c_{j}^{h,v}d_{n-j}^{h}$ and we conclude that $\xi^{-n-m}a_{n+m}^{h,v}\geq\xi^{-m}c_{m}^{h,v}\cdot\xi^{-n}d_{n}^{h}$ for every $n\in\mathbb{N}_{0}$. Hence
$\sum_{j=0}^{n}\xi^{-j-m}a_{j+m}^{h,v}\geq\xi^{-m}c_{m}^{h,v}\cdot\sum_{j=0}^{n}\xi^{-j}d_{j}^{h}\to+\infty$}
for $n\to\infty$, and $(A_{h,v})_{1/\xi}(1)_{1}=+\infty$. $\Box$
We proceed to $\xi$-convex weights.
@@
1. If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h,v})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}\neq 0$ and $(C_{h,v})_{1/\xi}(1)_{1}$ exist, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\big)\,.$}
2. If $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, and if $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}\neq 0$ and $C_{h,v}(1)_{1}$ exist, then
$A_{h,v}(1)_{1}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}\big)\,.$}
Proof. 1. By item 1 of Proposition 4.1 and by Proposition 2.31 we have $A_{h,v}(x)=C_{h,v}(x)D_{h}(x)=C_{h,v}(x)\frac{1}{1-\xi x}$. By Proposition 2.15,
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}\,.$
By item 2 of Proposition 4.1 we have $(C_{h,v})_{1/\xi}(x)^{2}=1-\frac{(B_{h,v})_{1/\xi}(x)}{(B_{h})_{1/\xi}(x)}$. By Proposition 2.22, $\big((C_{h,v})_{1/\xi}(1)_{1}\big)^{2}=1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}$. The stated formula follows.
2. Now, by item 1 of Proposition 4.1 and by Proposition 2.31, we have
$A_{h,v}(x)=C_{h,v}(x)\,.$
The rest of the argument is the same as in the previous case for $\xi=1$. $\Box$
Again, if we replace the assumption of non-vanishing of $B_{h,v}(1)_{1}$ with the existence of $E_{h,v}(1)_{1}$, we obtain the following variant of the theorem.
@@
1. If $\xi\in\mathbb{C}_{1}$, $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive $\xi$-convex light weight with $v\in V(h)$, and if $(B_{h,v})_{1/\xi}(1)_{1}$, $(B_{h})_{1/\xi}(1)_{1}$, $(C_{h,v})_{1/\xi}(1)_{1}$ and $(E_{h,v})_{1/\xi}(1)_{1}$ exist, then
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}\in\mathrm{sqrt}\big(1-\frac{(B_{h,v})_{1/\xi}(1)_{1}}{(B_{h})_{1/\xi}(1)_{1}}\big)\,$,}
where for $(B_{h})_{1/\xi}(1)_{1}=0$ we have $(B_{h,v})_{1/\xi}(1)_{1}=0$ and interpret the expression $\frac{0}{0}$ as $1$, so that then the limit is zero.
2. If $v\in\mathbb{N}\setminus\{1\}$, $h\colon\mathbb{N}_{2}\to\mathbb{C}$ is a $v$-transitive 0-convex light weight with $v\in V(h)$, and if $B_{h,v}(1)_{1}$, $B_{h}(1)_{1}$, $C_{h,v}(1)_{1}$ and $E_{h,v}(1)_{1}$ exist, then
$A_{h,v}(1)_{1}\in\mathrm{sqrt}\big(1-\frac{B_{h,v}(1)_{1}}{B_{h}(1)_{1}}\big)\,$,}
where for $B_{h}(1)_{1}=0$ we have $B_{h,v}(1)_{1}=0$ and interpret the $\frac{0}{0}$ as $1$, so that then $A_{h,v}(1)_{1}=0$.
Proof. 1. If $(B_{h})_{1/\xi}(1)_{1}\neq 0$, we use the previous proof. If $(B_{h})_{1/\xi}(1)_{1}=0$ then it follows from item 2 of Proposition 4.1 and Propositions 2.20 and 2.19 that also $(B_{h,v})_{1/\xi}(1)_{1}=0$. Item 3 of Proposition 4.1 and Proposition 2.20 give $(C_{h,v})_{\xi}(1)_{1}=0$. Thus, as in the previous theorem,
$\lim_{n\to\infty}\xi^{-n}a_{n}^{h,v}=(C_{h,v})_{1/\xi}(1)_{1}=0\,.$
2. Now, by item 1 of Proposition 4.1 and by Proposition 2.31, we have
$A_{h,v}(x)=C_{h,v}(x)\,.$
The rest of the argument is the same as in the previous case for $\xi=1$. $\Box$